# A Survey on Deep Learning for Theorem Proving

### v1

@@ 6.1 Challenges
Evaluation. Compared to traditional deep learning tasks such as classification, it is much harder to evaluate the performance of theorem provers comprehensively. First, results across different proof assistants are not comparable. Even though MiniF2F is available across multiple proof assistants, the impact due to proof automation (e.g., Isabelle has Sledgehammer (Böhme & Nipkow, 2010) whereas Lean does not) often outweighs the differences due to deep learning models. Second, the resource constraints (e.g., time, #attempts) during evaluation may impact the performance and the relative ranking of different methods. Compared to time constraints, #attempts favors bigger models such as LLMs. In contrast, a tight time constraint arising from real-time applications may favor simple but fast models. Without a specific application as the context, it is unclear what evaluation setting makes the most sense. We as a community still lack a systematic understanding of the trade-off, despite nascent efforts (Rute et al., 2024).
Human-AI interaction. One motivation for theorem proving is to assist human mathematicians (Castelvecchi, 2021; Sørensen et al., 2021). However, existing research has led to surprisingly few tools useful for them. Current methods are evaluated following standard deep learning protocols: running neural networks or querying LLMs in a Python program, testing the prover on a dataset, and calculating the percentage of successfully proved ones. This practice is misaligned with the needs of mathematicians. First, mathematicians need a tool that can be called easily in a proof assistant. Second, the tool must run on consumer CPUs with low latency. Third, instead of a performance measure, mathematicians care more about whether the prover can help with the specific theorems they are working on. These theorems are often out of the training distribution and pose a challenge for the prover to generalize to other domains. Although initial efforts have been made to develop user-oriented tools (Song et al., 2023; Welleck & Saha, 2023), there is abundant room in improving the user experience and exploring other forms of interaction, which requires close collaboration between deep learning researchers and mathematicians (Collins et al., 2023).
@@ 6.2 Future Directions
Combining deep learning, especially LLMs, with formal mathematics provides a promising avenue for enhancing the math capabilities of AI and may significantly impact various disciplines. We conclude our survey paper by listing a few future directions we are particularly excited about, envisioning significant strides in these burgeoning domains:
@@
Conjecturing. Beyond merely proving theorems, mathematicians would always explore theories in a domain, identify underlying problem structures, and formulate new conjectures. These explorative activities around conjecturing are indispensable for mathematicians but are currently limited in AI (Urban & Jakubüv, 2020; Johansson & Smallbone, 2023). By enabling AI to conjecture, it can explore the space of mathematics autonomously. Furthermore, when combined with theorem proving, such exploration can be used to discover new math knowledge. A direct application of conjecturing is to generate useful theorems and proofs, which could mitigate the data scarcity inherent in theorem proving.
@@
Verified code generation. As AI coding assistants such as GitHub Copilot become prevalent, it is increasingly important to be able to verify LLM-generated code. Proof assistants, especially Coq, have been widely used for software verification (Leroy et al., 2016). Therefore, methods for theorem proving surveyed in this paper may play a role in generating verified code. There is a plethora of problems to explore in this space. For example, one can prompt LLMs to synthesize inductive loop invariants (Kamath et al., 2023) or generate programs in verification-friendly languages such as Dafny (Sun et al., 2023).
@@
Math education. A roadblock to democratizing math education is the lack of qualified tutors to provide feedback to students (Liang et al., 2023b). Formal mathematics can potentially mitigate the problem by providing an environment for students to explore and receive automatic and reliable feedback. AI has demonstrated promise in guiding students in this process. For example, Kevin Buzzard reported that Lean Copilot could help prove a number of theorems in his undergraduate course and answer questions on Lean Zulip (Buzzard, 2024). To further integrate AI-driven formal tutoring into mainstream education, informalization is also essential to make these formal proofs accessible to students who are not familiar with formal languages. Looking ahead, we believe that LLMs and theorem proving could lead to intelligent tutors in a variety of STEM classes.

### v2

@@ 6.1 Challenges
Evaluation. Compared to traditional deep learning tasks such as classification, it is considerably more complex to evaluate the performance of theorem provers comprehensively. Firstly, results across different proof assistants are not directly comparable. Even though miniF2F is available across multiple proof assistants, the impact of proof automation tools (e.g., Isabelle has Sledgehammer (Böhme & Nipkow, 2010) whereas Lean does not) often outweighs the differences attributable to deep learning models. Secondly, resource constraints during evaluation, such as time and the number of attempts, can significantly affect performance and the relative ranking of different methods. While more attempts may favor larger models like LLMs, tight time constraints, especially in real-time applications, may favor simpler, faster models. Without a specific application as the context, it is unclear what evaluation setting makes the most sense. Thirdly, the use of pre-trained LLMs like GPT-4 may introduce data contamination issues (Rute et al., 2024), as these models might have been pre-trained on existing informal or formal testing problems, leading to unfair comparisons. We as a community still lack a systematic evaluation framework across different types of neural theorem provers, despite nascent efforts (Lamont et al., 2024; Rute et al., 2024).
Human-AI Interaction. One motivation for theorem proving is to assist human mathematicians (Castelvecchi, 2021; Sørensen et al., 2021). However, existing research has led to surprisingly few tools useful for them. Current methods are evaluated following standard deep learning protocols: running neural networks or querying LLMs in a Python program, testing the prover on a dataset, and calculating the percentage of successfully proved ones. This practice is misaligned with the needs of mathematicians. First, mathematicians need a tool that can be called easily in a proof assistant. Second, the tool must run on consumer CPUs with low latency. Third, instead of a performance measure, mathematicians care more about whether the prover can help with the specific theorems they are working on. These theorems are often out of the training distribution and pose a challenge for the prover to generalize to other domains. Although initial efforts have been made to develop user-oriented tools (Welleck & Saha, 2023; Song et al., 2024; Rute et al., 2024), there is abundant room to improve the user experience and explore other forms of interaction, which requires close collaboration between deep learning researchers and mathematicians (Collins et al., 2023).
@@ 6.2 Future Directions
Combining deep learning, especially LLMs, with theorem proving provides a promising avenue for enhancing the mathematical capabilities of AI and may significantly impact various disciplines. We conclude our survey paper by listing a few future directions we are particularly excited about, envisioning significant strides in these burgeoning domains:
@@
Conjecturing. Beyond merely proving theorems, mathematicians would always explore theories in a domain, identify underlying problem structures, and formulate new conjectures. These explorative activities around conjecturing are indispensable for mathematicians but are relatively limited in current deep learning approaches (Urban & Jakubüv, 2020; Johansson & Smallbone, 2023; Bengio & Malkin, 2024). By enabling AI to generate useful conjectures, it can explore the space of mathematics autonomously. When combined with theorem proving, such exploration can also be used to discover new mathematical knowledge. Furthermore, a direct application of conjecturing is to generate more theorem (and proof) data, which could mitigate the data scarcity inherent in theorem proving.
@@
Verified Code Generation. As AI coding assistants such as GitHub Copilot become prevalent, it is increasingly important to be able to verify LLM-generated code. Proof assistants, especially Coq, have been widely used for software verification (Leroy et al., 2016; Gu et al., 2016). Therefore, methods for theorem proving surveyed in this paper could potentially play a role in generating verified code. There is a plethora of problems to explore in this space. For example, one can train or prompt LLMs to synthesize program invariants (Pei et al., 2023; Kamath et al., 2023) or generate programs in verification-friendly languages such as Dafny (Sun et al., 2023), Verus (Yao et al., 2023), and F (Chakraborty et al., 2024).
@@
Math Education. A roadblock to democratizing math education is the lack of qualified tutors to provide feedback to students (Kumar et al., 2023). Formal mathematics can potentially mitigate this issue by providing an environment for students to explore and receive automatic and reliable feedback. AI has demonstrated promise in guiding students in this process. For example, Buzzard (2024) reported that Lean Copilot (Song et al., 2024) effectively assisted in proving a bunch of problems in his undergraduate course and solved questions on the Lean Zulip. To further integrate AI-driven formal tutoring into mainstream education, informalization is essential to make formal proofs accessible to students unfamiliar with formal languages. Looking ahead, we believe that theorem proving with LLMs will pave the way for intelligent tutors, enhancing math education for a broader audience.
Acknowledgments
We thank the anonymous reviewers for their insightful comments. This work was supported, in part, by Individual Discovery Grants from the Natural Sciences and Engineering Research Council of Canada, and the Canada CIFAR AI Chair Program.
Appendix A Lists of Datasets and State-of-the-Art Approaches
Dataset Language [^1] NL: natural language. DSL: domain-specific language. Size [^2] We use the number of tokens for pre-training datasets and the number of entries (e.g., definitions, premises, theorems) for other tasks. “” indicates the size of the testing/validation set. Source Task [^3] We list the main task of each dataset. AF: autoformalization. PS1: premise selection. PG: proofstep generation. PS2: proof search. PT: pre-training. TP: theorem proving (combination of premise selection, proofstep generation, and proof search).
NL-PS (Ferreira & Freitas, 2020a) NL 20,401 ProofWiki PS1
NaturalProofs (Welleck et al., 2021) NL 48,783 ProofWiki, Stacks, textbooks PS1
NaturalProofs-Gen (Welleck et al., 2022) NL $\sim$33K NaturalProofs PG
GamePad (Huang et al., 2019) Coq 1,602 Feit-Thompson theorem PG
CoqGym (Yang & Deng, 2019) Coq 70,856 123 open-source projects from Github TP
Proverbot9001 (Sanchez-Stern et al., 2020) Coq 501 CompCert TP
Graph2Tac (Rute et al., 2024) Coq $\sim$520K 120 open-source packages from opam TP
PISA (Jiang et al., 2021) Isabelle $\sim$183K Achieve of Formal Proofs, Isabelle standard library TP
MAPL (Mikuła et al., 2024) Isabelle $\sim$433K Achieve of Formal Proofs, Isabelle standard library PS1
Selene (Zhang et al., 2024) Isabelle 360 seL4 TP
FVELer (Lin et al., 2024b) Isabelle 29,125 seL4 TP
LeanStep (Han et al., 2022) Lean $\sim$20K Lean core library, mathlib TP
TRIGO (Xiong et al., 2023) Lean 427 trigonometry problems from tiku TP
LeanDojo (Yang et al., 2023) Lean 98,734 mathlib TP
MLFMF (Bauer et al., 2023) Lean, Agda 270,647 mathlib, Agda standard library, UniMath, TypeTopology PS1
miniCodeProps (Lohn & Welleck, 2024) Lean, Haskell 177* Tons of Inductive Programs TP
LEAN-GitHub (Wu et al., 2024) Lean 28,597 147 repositories on the web TP
FIMO (Liu et al., 2023) NL, Lean 149 IMO Shortlisted Problems AF
ProofNet (Azerbayev et al., 2023) NL, Lean 371 undergraduate textbooks AF
miniF2F (Zheng et al., 2022) NL, Lean, Isabelle, Metamath, HOL Light 488 IMO, AIME, AMC, MATH, custom TP
PutnamBench (Tsoukalas et al., 2024) NL, Lean, Isabelle 640 William Lowell Putnam Mathematical Competition TP
HolStep (Kaliszyk et al., 2017) HOL Light 11,410 multivariate analysis library & Kepler conjecture PS1
HOList (Bansal et al., 2019) HOL Light 31,662 Kepler conjecture TP
MPTP2078 (Alama et al., 2014) Mizar, TPTP 2,078 Mizar Mathematical Library PS1, PS2
Mizar40 (Kaliszyk & Urban, 2015b) Mizar, TPTP $\sim$58k Mizar Mathematical Library PS1, PS2
M2K (Kaliszyk et al., 2018) Mizar, TPTP 2,003 Mizar40 PS2
TPTP (Sutcliffe, 2017) TPTP 11,395 ATP system evaluation PS2
TacticToe (Gauthier et al., 2020) HOL4 7,164 HOL4 standard library TP
set.mm (Lample et al., 2022) Metamath 37,091 set.mm TP
UniGeo (Chen et al., 2022) NL, DSL 9,543 high school geometry problems from IXL PG
IMO-AG-30 (Trinh et al., 2024) DSL 30 IMO geometry problems TP
INT (Wu et al., 2021b) DSL 1,501K synthetic inequality problems TP
AlphaGeometry (Trinh et al., 2024) DSL $\sim$100M synthetic geometry problems TP
MMA (Jiang et al., 2023a) NL, Lean, Isabelle 332,774 autoinformalization from Achieve of Formal Proofs and mathlib AF
FormL4 (Lu et al., 2024) NL, Lean 17,461 autoinformalization from mathlib, Arithmo test set AF
MUSTARD (Huang et al., 2024) NL, Lean 5,866 autoformalization from synthetic problems TP
DeepSeek-Prover (Xin et al., 2024a) NL, Lean 8,066,621 autoformalization from competition problems TP
Lean Workbook (Ying et al., 2024a) NL, Lean $\sim$57K autoformalization from Art of Problem Solving AF
WebMath (Polu & Sutskever, 2020) mixed 35B GitHub, arXiv, Stack Exchange PT
Proof-Pile (Azerbayev et al., 2023) mixed 8.3B arXiv, Stack Exchange, formal libraries, ProofWiki, Wikipedia, books, MATH PT
MathPile (Wang et al., 2023b) mixed 9.5B arXiv, textbooks, Stack Exchange, Wikipedia, ProofWiki, Web PT
OpenWebMath (Paster et al., 2024) mixed 14.7B forum posts, educational content, reference pages, scientific papers, blogs, and others PT
Proof-Pile-v2 (Azerbayev et al., 2024) mixed 55B Github, Stack, formal proof steps, OpenWebMath, arXiv PT
DeepSeekMath (Shao et al., 2024) mixed 120B OpenWebMath, Web PT
InternLM-Math (Ying et al., 2024b) mixed 125B Knowledge Pile, Proof-Pile-v2, synthetic data PT
Table 1: Summary of existing datasets for theorem proving. For datasets derived from generation methods, we primarily highlight several key examples here.
Language Dataset Best Accuracy (%) Evaluation Metric Method
Coq CoqGym 33.8 timeout of 600s Diva (First & Brun, 2022) + CoqHammer (Czajka & Kaliszyk, 2018)
Lean LeanDojo (random) 57.7 pass@1 + timeout of 600s temperature scaling (Gloeckle et al., 2023)
ProofNet (valid/test) 49.2/53.2 top-1 accuracy@50 type checking (Poiroux et al., 2024)
miniF2F (valid/test) 60.2/46.3 cumulative/pass@64 DeepSeek-Prover (Xin et al., 2024a)
Isabelle miniF2F (valid/test) 57.0/51.2 pass@100/pass@200 LEGO-Prover (Xin et al., 2024b)/Lyra (Zheng et al., 2024)
PISA 71.0 pass@300 + timeout of 500s Magnushammer (Mikuła et al., 2024)+Thor(Jiang et al., 2022)
Metamath set.mm (valid/test) 81.2/72.4 pass@32 HTPS (Lample et al., 2022)
Mizar MPTP2078 75.5 timeout of 100s NIAGRA (Fokoue et al., 2023)
Table 2: Summary of state-of-the-art methods and their performance on several key datasets.